# CryptoMamba: Leveraging State Space Models for Accurate Bitcoin Price Prediction

### v1

@@ VI Application in Real-World Trading
These results emphasize the practical utility of CryptoMamba for real-world financial applications. By combining accurate predictions with trading strategies, CryptoMamba demonstrates its ability to excel under diverse market conditions, making it a reliable solution for adaptive financial forecasting.
VII Conclusion and Future Work
@@
In this work, we introduced CryptoMamba, a novel Mamba-based architecture tailored for financial time-series forecasting, and demonstrated its effectiveness in predicting Bitcoin prices. Leveraging state space models (SSMs) with Mamba enhancements, CryptoMamba efficiently captures long-range dependencies and outperforms traditional models such as LSTM, Bi-LSTM, GRU, and S-Mamba in both predictive accuracy and financial performance. Our results show that CryptoMamba, especially in its volume-inclusive variant, achieves the highest returns in real-world trading scenarios, highlighting the importance of incorporating volume data for enhanced market insights.
@@
The trading simulations using Vanilla and Smart algorithms emphasize the practical utility of CryptoMamba in real-world trading. By employing trading algorithms as evaluation tools, we were able to effectively compare prediction models in terms of their financial impact, moving beyond regression metrics alone. CryptoMamba design not only proves its capability in modeling cryptocurrency markets but also showcases its potential for forecasting other time series data, such as stocks or commodities, where long-range dependencies are critical.

### v2

@@ 6 Application in Real-World Trading
These results emphasize the practical utility of CryptoMamba for real-world financial applications. By combining accurate predictions with trading strategies, CryptoMamba demonstrates its reliability and adaptability across diverse market conditions for financial forecasting.
7 Conclusion and Future Work
@@
In this work, we introduced CryptoMamba, a novel Mamba-based architecture tailored for financial time-series forecasting, and demonstrated its effectiveness in predicting Bitcoin prices. Leveraging state space models (SSMs) with Mamba enhancements, CryptoMamba efficiently captures long-range dependencies and outperforms traditional models such as LSTM, Bi-LSTM, GRU, iTransformer, and S-Mamba in both predictive accuracy and financial performance. Our results show that CryptoMamba, especially in its volume-inclusive variant, achieves the highest returns in real-world trading scenarios, highlighting the importance of incorporating volume data for enhanced market insights.
@@
The trading simulations using Vanilla, Smart, and Extended Smart algorithms emphasize the practical utility of CryptoMamba in real-world trading. By employing trading algorithms as evaluation tools, we were able to effectively compare prediction models in terms of their financial impact, moving beyond regression metrics alone. These results highlight the versatility of CryptoMamba in adapting to different market conditions and underline its ability to achieve profitability even under highly volatile scenarios. CryptoMamba design not only proves its capability in modeling cryptocurrency markets but also showcases its potential for forecasting other time series data, such as stocks or commodities, where long-range dependencies are critical.
@@
Appendix
Appendix A Metrics
Regression Metrics.
For evaluating the forecasting accuracy of our models, we use three standard metrics: Root Mean Squared Error (RMSE), Mean Absolute Percentage Error (MAPE), and Mean Absolute Error (MAE). The formulas for these metrics are presented in Table 5, where $y_{i}$ represents the actual value, $\hat{y}_{i}$ is the predicted value, and $n$ is the total number of predictions.
Table 5: Definitions of the evaluation metrics
Metric Formula
RMSE $\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$
MAPE $\frac{100}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|$
MAE $\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$
Maximum Drawdown (MDD).
Maximum Drawdown (MDD) quantifies the largest peak-to-trough decline in a portfolio’s value over a given time period. It measures the worst possible loss an investor could have faced if buying at a local maximum and selling at the subsequent minimum. MDD is defined as:
$\text{MDD}=\max_{t\in[1,T]}\left(\frac{\max_{i\in[1,t]}P_{i}-P_{t}}{\max_{i\in[1,t]}P_{i}}\right)$ (1)
where:
  • $P_{t}$ is the portfolio value at time step $t$,
  • $T$ is the total number of time steps,
  • $\max_{i\in[1,t]}P_{i}$ is the historical peak value up to time $t$.
This metric expresses the largest percentage drop from a historical peak, helping assess downside risk and the volatility of a trading strategy. Lower MDD values indicate better risk control.
Appendix B Extended Smart Trading Algorithm
The Extended Smart Trading Algorithm builds upon the Smart strategy by incorporating short-selling capabilities, enabling the model to profit even during market downturns. Like the Smart algorithm, it defines a risk-based interval $[y_{\min},y_{\max}]$ around the predicted price $y$ and adjusts the trading action based on the current price $x$ relative to this interval. However, in scenarios where the predicted price is lower than today’s price, the algorithm can go short (i.e., sell more than the currently held bitcoin). To prevent excessive risk exposure, we introduce a maximum short position limit, restricting the amount of negative bitcoin the strategy can hold to 0.002 BTC. This constraint ensures controlled risk while enabling more aggressive trading decisions that can outperform traditional long-only strategies in volatile or bearish markets. The detailed steps of this algorithm are provided in Algorithm 3.
Algorithm 3 Extended Smart Trading Algorithm
0:  Prediction, Today’s Price, risk, maximum negative bitcoin
[…]
2:  $y\leftarrow\text{Prediction}$
3:  $y_{\max}\leftarrow\left(1+\frac{\text{risk}}{100}\right)y$
4:  $y_{\min}\leftarrow\left(1-\frac{\text{risk}}{100}\right)y$
5:  if $x\geq y$ then
6:     if $x\geq y_{\max}$ then
7:        Sell as much as shares to reach the maximum short position
8:     else
9:        if you have positive shares then
10:           Sell $\frac{x-y}{y_{\max}-y}$ of your shares
11:        end if
12:     end if
13:  else
14:     if $x\leq y_{\min}$ then
15:        Buy with all of your money
16:     else
17:        Buy with $\frac{y-x}{y-y_{\min}}$ of your money
18:     end if
19:  end if